the force of gravity on the surface of the moon is about 1/6 of the force of gravity on the earth's surface. if my vertical jump is 2.3 how high could i jump on the surface of the moon?
Answer by khwang(438)

You should have units in physics questions.
Note, by the formula about accereation v^2 = 1/2 gH. You jump at the
same initial velocity either on the earth or on the moon.
Hence,the max height you can jump on a planet is inverse proportional
to g(the gravitional acc. constant )

So,the height you can jump on the moon is 6*2.3 = 13.8 meters (or ft ???)


Kenny